A tube of mercury at a room temperature of 22.4 C has a volume of 10.6 mL between the sealed end of the tube and the mercury. The sun rises and shines through a window on the tube and warms it to 27.8 C. If the atmospheric pressure remains constant, what is the new volume between the sealed end of the tube and the mercury?



1
Expert's answer
2022-03-29T16:22:04-0400

V1/T1 = V2/T2

V1 = 10.6 mL

T1 = 22.4 + 273.15 = 295.55 K

T2 = 27.8 + 273.15 = 300.95 K

V2 = V1T2 / T1 = 10.6 * 300.95 / 295.55 = 10.8 (mL)